Find the resistance of a circuit runs with 20 A and 54.2 nV



1
Expert's answer
2022-05-24T13:56:20-0400
"R=\frac{V}{I}=\frac{54.2*10^{-9}}{20}=2.71*10^{-9}\:\Omega"